Logiweb bibliography of check

Up Help

[0] check

[1] base

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-06-21.UTC:16:52:09.826502 = MJD-53907.TAI:16:52:42.826502 = LGT-4657625562826502e-6